# A Closer Look at Model Collapse: From a Generalization-to-Memorization Perspective

### v2

@@ A.2 Generalization and Memorization
Recent studies [Yoon et al., 2023, Zhang et al., 2024] have identified two distinct learning regimes in diffusion models, depending on the size of the training dataset and the model’s capacity: (1) Memorization regime, when models with sufficient capacity are trained on limited datasets, they tend to memorize the training data; and (2) Generalization regime, as the number of training samples increases, the model begins to approximate the underlying data distribution and generate novel samples. To investigate the transition between these regimes, Wang et al. [2024b] show that the number of training samples required for the transition from memorization to generalization scales linearly with the intrinsic dimension of the dataset. In addition, the analysis of training and generation accuracies in Wang et al. [2024c] provides a potential step toward quantifying generalization. Zhang et al. [2025a] propose a theoretically grounded and computationally efficient metric, Probability Flow Distance (PFD), to measure the generalization ability of diffusion models. Specifically, PFD quantifies the distance between distributions by comparing their noise-to-data mappings induced by the probability flow ODE. Meanwhile, concurrent work also explores memorization and generalization separately. To understand generalization, studies such as Kadkhodaie et al. [2024], An et al. [2024] attribute the generalization to implicit bias introduced by network architectures. Other works study the generalized distribution using Gaussian models [Wang and Vastola, 2024, Li et al., 2024] and patch-wise optimal score functions [Niedoba et al., 2024, Kamb and Ganguli, 2024]. As for memorization, it is investigated in both unconditional and conditional [Gu et al., 2025], as well as the text-to-image diffusion models [Carlini et al., 2023, Somepalli et al., 2023a]. Additionally, methods to mitigate memorization in diffusion models have been proposed in Somepalli et al. [2023b], Wen et al. [2024]. Distinct from prior work, our study is the first to establish a connection between model collapse and the transition from generalization to memorization. This connection not only offers a novel perspective to understand model collapse but also provides insights to mitigate it by mitigating memorization.
@@ A.3 Data Selection
Sorscher et al. [2022], Kolossov et al. [2024], Hemmat et al. [2025] focus on data pruning techniques, but not necessarily in the context of self-consuming loops. While both their approaches and ours demonstrate the benefits of data selection, there are several key differences:
  • Objective and evaluation: Sorscher et al. [2022], Kolossov et al. [2024], Hemmat et al. [2025] primarily study how pruning improves scaling laws, achieving higher accuracy as dataset size varies. In contrast, our work examines how model performance evolves over an iterative training loop with a fixed dataset size, focusing on the generalization-to-memorization transition.
@@
  • Pruning strategy: Methods in Sorscher et al. [2022], Kolossov et al. [2024], Hemmat et al. [2025] largely rely on per-sample evaluation, whereas our approach considers global dataset structure and relationships between samples. While this may lead to increased computational complexity, it opens up new possibilities for designing pruning criteria beyond per-sample evaluation.
  • Entropy definition: In Hemmat et al. [2025], while the authors use a generative model to sample images, the goal is to improve the classification performance of a classifier. And the entropy used in their method is defined in the prediction space of a classifier, which is different from the entropy of a dataset measured in our work.
@@ B.1 Network Structure
For CIFAR-10 and FFHQ, we use a UNet-based backbone, taking RGB images as inputs and predicting noise residuals. Our implementation is based on the Hugging Face Diffusers base code [von Platen et al., 2022]. The architecture hyperparameters of the neural network are listed as follows:
@@ B.2 Implementation Details
In this paper, we use DDPM as our generative method. For efficiency, we use the FP-16 mixing precision to train our models, which is inherently implemented by the Hugging Face Diffusers codebase. The batch size of training all datasets is set to be $128$. A $1000$-step denoising process is used as the reverse process. For CIFAR-10, the epoch number is set to be $500$; for FFHQ, the epoch number is set to be $1000$. We use the Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. Other experimental hyperparameters are exactly the default values in the original Hugging Face Diffusers codebase. We use the DINOv2 model [Oquab et al., 2024] to extract features of images and then calculate the distance between two images in the feature space. We use the InceptionV3 model [Szegedy et al., 2016] to extract features of images to calculate the FID score. All experiments are conducted on a single NVIDIA A-100 GPU.
@@ C.1 Training on More Samples
  • Incorporating more data into the training-sampling recursion can indeed mitigate the rate of model collapse. Compared to the vanilla replacement paradigm (i.e., the first line), using $2N$ images (second line) yields improved performance. This aligns with prior findings [Kazdan et al., 2024, Fu et al., 2024] that sample size is a key factor influencing the collapse rate.
@@ D.2 The Robust Relationship between Generalization Score and Entropy
In Figure˜2, we plot the results on CIFAR-10 datasets, demonstrating the dataset size-independent relationship between generalization score and estimated entropy. This subsection further incorporates the results from FFHQ and the accumulate paradigm.
@@ D.4 Generated Image Distribution Becomes Spiky
To quantitatively measure the spiky degree of the empirical image distribution, we further adopt the Mean Nearest Neighbor Distance (MNND) [Jain et al., 2004, Abbar et al., 2013], which removes the data-independent constants and logarithm in the KL estimator:
$\mathtt{MNND}(\mathcal{D}_{t})\triangleq\mathtt{Dist}(\mathcal{D}_{t},\mathcal{D}_{t})=\frac{1}{|{\mathcal{D}_{t}}|}\sum_{\bm{x}\in{\mathcal{D}_{t}}}\min_{\bm{z}\in{\mathcal{D}_{t}}\backslash\bm{x}}d(\bm{x},\bm{z}).$ (3)
@@
$e^{\frac{\hat{H}_{1}(\mathcal{D}_{t})-B}{d}}\leq\text{MNND}(\mathcal{D}_{t}),$ (4)
where $\hat{H}_{1}(\mathcal{D}_{t})$ represents the estimated entropy of $\mathcal{D}_{t}$ with $k=1$ and $B$ is a constant offset given a fixed dataset size.
We want to clarify the nuance between MNND and variance. A spiky distribution does not imply that all data points are concentrated in a single small region—this behavior has already been illustrated by the variance collapse behavior [Shumailov et al., 2024, Kazdan et al., 2024, Bertrand et al., 2024]. Specifically, Figure˜20 shows that the variance of the generated dataset only slightly decreases along the successive iterations and is far from complete collapse. On the contrary, MNND decreases almost exponentially and reaches a small value after $10$ iterations. Thus, the results align with the prior claim in Schaeffer et al. [2025], Kazdan et al. [2024] that the collapse of variance progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the collapse in MNND emerges at an early stage, highlighting a critical memorization issue caused by training on synthetic data.
@@ D.5 Self-consuming Loop with Fresh New Data
Alemohammad et al. [2024a] show that incorporating fresh real data can further mitigate model collapse. In this section, we conduct experiments to verify the effectiveness of our methods in this paradigm. The candidate data pool in each iteration is jointly composed of three sources: (1) the original real data, (2) synthetic data generated by the previous model, and (3) fresh real data that was not used in earlier iterations. This experiment is designed to simulate real-world deployment, where generative models are continuously updated with a mixture of prior synthetic data and incoming fresh real data, and the training budget also increases. Concretely, we first train Model A on 32,768 real images, which then generates 10,000 synthetic images. We construct a data pool—simulating an Internet-scale source—by combining the 32,768 original real images, 10,000 synthetic images, and 10,000 additional fresh real images. From this pool, our entropy-based selection method chooses 40,000 images as the training dataset for Model C. For comparison, Model B is trained on 40,000 randomly selected images from the same pool.
@@ D.6 Additional Metric
Since our Greedy Section method is performed on the feature space extracted by the DINO model, we also use FDDINO [Stein et al., 2023] as an alternate for FID to measure the quality of the generated images. As we show in Figure˜21, the Greedy Selection method can also improve the FDDINO metric compared to the vanilla paradigms.
@@
D.7 Visualization of the Improvement on Generalization Score

### v3

@@ A.2  Generalization and Memorization
Recent studies [yoon2023diffusion, pmlr-v235-zhang24cn] have identified two distinct learning regimes in diffusion models, depending on the size of the training dataset and the model’s capacity: (1) Memorization regime, when models with sufficient capacity are trained on limited datasets, they tend to memorize the training data; and (2) Generalization regime, as the number of training samples increases, the model begins to approximate the underlying data distribution and generate novel samples. To investigate the transition between these regimes, [wang2024diffusion] show that the number of training samples required for the transition from memorization to generalization scales linearly with the intrinsic dimension of the dataset. In addition, the analysis of training and generation accuracies in [wang2024evaluating] provides a potential step toward quantifying generalization. [DBLP:journals/corr/abs-2505-20123] propose a theoretically grounded and computationally efficient metric, Probability Flow Distance (PFD), to measure the generalization ability of diffusion models. Specifically, PFD quantifies the distance between distributions by comparing their noise-to-data mappings induced by the probability flow ODE. Meanwhile, concurrent work also explores memorization and generalization separately. To understand generalization, studies such as [kadkhodaie2023generalization, an2024inductive] attribute the generalization to implicit bias introduced by network architectures. Other works study the generalized distribution using Gaussian models [wang2024unreasonable, li2024understanding] and patch-wise optimal score functions [niedoba2024towards, kamb2024analytic]. As for memorization, it is investigated in both unconditional and conditional [DBLP:journals/tmlr/GuDPLL025], as well as the text-to-image diffusion models [Carlini2023Extracting, somepalli2023diffusion]. Additionally, methods to mitigate memorization in diffusion models have been proposed in [somepalli2023understanding, wen2024detecting]. Distinct from prior work, our study is the first to establish a connection between model collapse and the transition from generalization to memorization. This connection not only offers a novel perspective to understand model collapse but also provides insights to mitigate it by mitigating memorization.
@@ A.3  Data Selection
[DBLP:conf/nips/SorscherGSGM22, DBLP:conf/iclr/KolossovMT24, DBLP:journals/corr/abs-2502-15588] focus on data pruning techniques, but not necessarily in the context of self-consuming loops. While both their approaches and ours demonstrate the benefits of data selection, there are several key differences:
  • Objective and evaluation: [DBLP:conf/nips/SorscherGSGM22, DBLP:conf/iclr/KolossovMT24, DBLP:journals/corr/abs-2502-15588] primarily study how pruning improves scaling laws, achieving higher accuracy as dataset size varies. In contrast, our work examines how model performance evolves over an iterative training loop with a fixed dataset size, focusing on the generalization-to-memorization transition.
@@
  • Pruning strategy: Methods in [DBLP:conf/nips/SorscherGSGM22, DBLP:conf/iclr/KolossovMT24, DBLP:journals/corr/abs-2502-15588] largely rely on per-sample evaluation, whereas our approach considers global dataset structure and relationships between samples. While this may lead to increased computational complexity, it opens up new possibilities for designing pruning criteria beyond per-sample evaluation.
  • Entropy definition: In [DBLP:journals/corr/abs-2502-15588], while the authors use a generative model to sample images, the goal is to improve the classification performance of a classifier. And the entropy used in their method is defined in the prediction space of a classifier, which is different from the entropy of a dataset measured in our work.
@@ B.1  Network Structure
For CIFAR-10 and FFHQ, we use a UNet-based backbone, taking RGB images as inputs and predicting noise residuals. Our implementation is based on the Hugging Face Diffusers base code [von-platen-etal-2022-diffusers]. The architecture hyperparameters of the neural network are listed as follows:
@@ B.2  Implementation Details
In this paper, we use DDPM as our generative method. For efficiency, we use the FP-16 mixing precision to train our models, which is inherently implemented by the Hugging Face Diffusers codebase. The batch size of training all datasets is set to be $128$. A $1000$-step denoising process is used as the reverse process. For CIFAR-10, the epoch number is set to be $500$; for FFHQ, the epoch number is set to be $1000$. We use the Adam optimizer with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. Other experimental hyperparameters are exactly the default values in the original Hugging Face Diffusers codebase. We use the DINOv2 model [DBLP:journals/tmlr/OquabDMVSKFHMEA24] to extract features of images and then calculate the distance between two images in the feature space. We use the InceptionV3 model [DBLP:conf/cvpr/SzegedyVISW16] to extract features of images to calculate the FID score. All experiments are conducted on a single NVIDIA A-100 GPU.
@@ C.1  Training on More Samples
  • Incorporating more data into the training-sampling recursion can indeed mitigate the rate of model collapse. Compared to the vanilla replacement paradigm (i.e., the first line), using $2N$ images (second line) yields improved performance. This aligns with prior findings [DBLP:journals/corr/abs-2410-16713, DBLP:conf/icml/Fu [^0] T24] that sample size is a key factor influencing the collapse rate.
@@ D.2  The Robust Relationship between Generalization Score and Entropy
In Figure˜2, we plot the results on CIFAR-10 datasets, demonstrating the dataset size-independent relationship between generalization score and estimated entropy. This subsection further incorporates the results from FFHQ and the accumulate paradigm to show a more general relation.
@@ D.4  Generated Image Distribution Becomes Spiky
To quantitatively measure the spiky degree of the empirical image distribution, we further adopt the Mean Nearest Neighbor Distance (MNND) [DBLP:conf/pakdd/JainSH04, DBLP:conf/compgeom/AbbarAIMV13], which removes the data-independent constants and logarithm in the KL estimator:
[…]
@@
[…]
We want to clarify the nuance between MNND and variance. A spiky distribution does not imply that all data points are concentrated in a single small region—this behavior has already been illustrated by the variance collapse behavior [DBLP:journals/nature/ShumailovSZPAG24, DBLP:journals/corr/abs-2410-16713, DBLP:conf/iclr/BertrandBDJG24]. Specifically, Figure˜20 shows that the variance of the generated dataset only slightly decreases along the successive iterations and is far from complete collapse. On the contrary, MNND decreases almost exponentially and reaches a small value after $10$ iterations. Thus, the results align with the prior claim in [DBLP:journals/corr/abs-2503-03150, DBLP:journals/corr/abs-2410-16713] that the collapse of variance progresses at such a slow pace that it is rarely a practical concern in real-world applications. In contrast, the collapse in MNND emerges at an early stage, highlighting a critical memorization issue caused by training on synthetic data.
@@ D.5  Self-consuming Loop with Fresh New Data
[DBLP:conf/iclr/AlemohammadCLHB24] show that incorporating fresh real data can further mitigate model collapse. In this section, we conduct experiments to verify the effectiveness of our methods in this paradigm. The candidate data pool in each iteration is jointly composed of three sources: (1) the original real data, (2) synthetic data generated by the previous model, and (3) fresh real data that was not used in earlier iterations. This experiment is designed to simulate real-world deployment, where generative models are continuously updated with a mixture of prior synthetic data and incoming fresh real data, and the training budget also increases. Concretely, we first train Model A on 32,768 real images, which then generates 10,000 synthetic images. We construct a data pool—simulating an Internet-scale source—by combining the 32,768 original real images, 10,000 synthetic images, and 10,000 additional fresh real images. From this pool, our entropy-based selection method chooses 40,000 images as the training dataset for Model C. For comparison, Model B is trained on 40,000 randomly selected images from the same pool.
@@ D.6  Additional Metric
Since our Greedy Section method is performed on the feature space extracted by the DINO model, we also use FDDINO [DBLP:conf/nips/SteinCHSRVLCTL23] as an alternate for FID to measure the quality of the generated images. As we show in Figure˜21, the Greedy Selection method can also improve the FDDINO metric compared to the vanilla paradigms.
@@
D.7  Additional Results for Section˜5